# MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning

### v1

@@ 1  Step 4: Initial Embedding construction:
4If Encoding Method 1: Concatenate $\psi_{{f}_{r}}(s_{j},:)$, $j=1,...K$, $r=1,...,D$ in a matrix form $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$.
If Encoding Method 2: Concatenate $\psi_{{f}_{r}}(s_{j},:)$ for $r=1,...,D$ in a matrix form $\psi_{\mathbf{x}}(s_{j},:,:)$ for each fixed scale $s_{j}$, representing all scales using 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
@@
5.2 Optimization
@@
We propose to use optimization based on SGD that begins with the initial embedding within the spectral graph domain. While employing positive and negative sampling, our aim is to optimize the embedding space directly in the SGW domain, revealing the intrinsic structure of manifold data while retaining high-frequency information associated with the graph Laplacian. This optimization within the SGW domain provides more effective regularization, allowing for the direct removal of noise from each spectral band. We outline the steps of the proposed approach, presenting several alternatives for optimization through the incorporation of multi-scale representations (refer to method 1 and method 2 in the algorithm description below).
@@ 1. Optimize Embedding Method 1:
$\frac{\partial\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})}{\partial\psi_{\mathbf{x}_{i}}}\sim\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}}(\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}})-\sum_{j}\frac{1}{||\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}}||^{2}}v_{ij}^{\psi_{\mathbf{x}}}(\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}})$ (6)
Our optimization process involves using SGD with positive and negative sampling, similar to recent graph embeddings method such as UMAP and LargeVis [26]. Positive edge samples are associated with attraction (first term on the right hand side), while negative samples (second term on the right hand side) refer to a pair of nodes that are not connected in the graph, which create repulsion among dissimilar points.
Output: Regularized embedding space $\tilde{\psi}_{\mathbf{x}}$.
2. 3D-Tensor Based Optimization: Embedding Method 2:
In method 2, the encoded manifold representation given is a 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$. A SGD update rule at iteration $t$ applied to the 3D tensor is:
$(\tilde{\psi}_{\mathbf{x}}^{(t+1)})_{i,j,k}=(\tilde{\psi}_{\mathbf{x}}^{(t)})_{i,j,k}-\alpha\frac{\partial\mathcal{L}}{(\partial\psi_{\mathbf{x}})_{i,j,k}}$ (7)
where $\alpha$ is the learning parameter. In our case we employ the cross entropy loss function:
$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})=\sum_{i,j,k}\left(w_{ij}\mbox{log}\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}+(1-w_{ij})\mbox{log}\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right)$ (8)
We apply the optimization based on SGD with respect to each scale $s_{k}$:
$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)|\mathbf{W})=\sum_{i,j}\left(w_{ij}\mbox{log}\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}+(1-w_{ij})\mbox{log}\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right)$ (9)
which has the gradient of the loss is approximated by:
$\frac{\partial\mathcal{L}(\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)|\mathbf{W})}{\partial\psi_{\mathbf{x}_{i}}(s_{k},:,:)}=\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}(\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:))$
$-\sum_{j}\frac{1}{||\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:)||^{2}}v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}(\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:))$
$\tilde{\psi}_{\mathbf{x}}=\sum_{k}\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)$ (10)
In the 3D tensor based optimization method 2, the dimensionality of the final embedding $\tilde{\psi}_{\mathbf{x}}\in\mathbb{R}^{D}\times\mathbb{R}^{N}$, maintaining the same dimensionality as the original input, and thus there exists a one-to-one correspondence between the coordinates of the input features $\mathbf{x}$ and $\tilde{\psi}_{\mathbf{x}}$. This correspondence ensures that each dimension of the feature space is directly mapped to a single dimension in the embedding space $\tilde{\psi}_{\mathbf{x}}$.
Sampling approaches from network features: We propose a strategic sampling of edges, which further refines the optimization by assessing the significance of edges with respect to the topological structure of the graph. Specifically, we construct probability distribution over $V$ (or $\mathbf{W}$), providing the significance of each node/edge, and sample from this distribution. We also propose a novel method to implement the sampling strategy, rather than selecting edges randomly as in [21, 26]. Refer to Appendix for more details.
In this section, we characterize the theoretical properties of the representation power of the SGW operator by considering functions sampled from the Paley-Wiener spaces [22] on combinatorial graphs. Pesenson introduced the Paley-Wiener spaces and demonstrated that Paley-Wiener functions of low type are uniquely determined by their values on certain subsets, known as uniqueness sets $U$. We show that the SGW operator can represent functions $f$ within the Paley-Wiener space more efficiently than the graph Laplacian operator $\mathcal{L}$. This efficiency is demonstrated by showing that the SGW operator is more effective in representing functions with larger bandwidth $\omega$ in the Paley-Wiener spaces (i.e., with higher frequencies) using subsets of nodes from the uniqueness sets $U$.
The $\Lambda$-set is defined as follows: a set of vertices $S\subset V$ is a $\Lambda(S)$-set if any $\varphi\in L_{2}(S)$ satisfies the inequality $||\varphi||\leq\Lambda||\mathcal{L}\varphi||$ for some constant $\Lambda(S)>0$. The infimum of all $\Lambda>0$ for which $S$ is a $\Lambda$-set is denoted by $\Lambda$.
Theorem 1
The proof is provided in the Appendix.
7 Experimental Results
We demonstrates the advantage of our approach over existing approaches using both synthetic and real datasets. We evaluate our method by testing its clustering performance using the output embedding and comparing it to several representative methods, including UMAP, t-SNE, Diffusion Maps, ISOMAP, and HeatGeo [19]. For HeatGeo, we hyperparameter tune the performance on each dataset and report the best one in Section 11. For each dataset, we briefly describe their properties and leave fuller details in Section 9.

### v4

@@ 1  Step 4: Initial Embedding construction:
  Step 5: Concatenate $\psi_{{f}_{r}}(s_{j},:)$ for $r=1,...,D$ in a matrix form $\psi_{\mathbf{x}}(s_{j},:,:)$ for each fixed scale $s_{j}$, representing all scales using 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
@@
3.2 Step 2: Optimization design using multi-scale network structure.
@@
We propose to use optimization based on SGD that begins with the initial embedding within the spectral graph domain. While employing positive and negative sampling, our aim is to optimize the embedding space directly in the SGW domain, revealing the intrinsic structure of manifold data while retaining high-frequency information associated with the graph Laplacian. This optimization within the SGW domain provides more effective regularization, allowing for the direct removal of noise from each spectral band. We outline the steps of the proposed approach.
3D-Tensor Based Optimization for graph Embedding
The encoded manifold representation is a 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$. An SGD update rule at iteration $t$, applied to the 3D tensor, is:
$(\tilde{\psi}_{\mathbf{x}}^{(t+1)})_{i,j,k}=(\tilde{\psi}_{\mathbf{x}}^{(t)})_{i,j,k}-\alpha\frac{\partial\mathcal{L}}{\partial(\psi_{\mathbf{x}})_{i,j,k}}$ (2)
where $\alpha$ is the learning parameter. In this case, we employ the cross-entropy loss function:
$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})=\sum_{i,j,k}$ $\left(w_{ij}\log\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right.$
$\left.+(1-w_{ij})\log\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right)$ (3)
where $v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}=\frac{1}{1+||\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:)||^{2}}$. We apply the optimization based on SGD with respect to each scale $s_{k}$, where the gradient of the loss is approximated by:
$\frac{\partial\mathcal{L}\left(\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)\mid\mathbf{W}\right)}{\partial\psi_{\mathbf{x}_{i}}(s_{k},:,:)}$ $=\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}\,r_{ij}(s_{k})$ (4)
$\quad-\sum_{j}\frac{1}{\|r_{ij}(s_{k})\|^{2}}v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}\,r_{ij}(s_{k})$ (5)
where $r_{ij}(s_{k})=\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:)$.
$\tilde{\psi}_{\mathbf{x}}=\sum_{k}\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)$ (6)
Note that the dimensionality of the final embedding
$\tilde{\psi}_{\mathbf{x}}\in\mathbb{R}^{D}\times\mathbb{R}^{N}$, maintaining the same dimensionality as the original input, and thus there exists a one-to-one correspondence between the coordinates of the input features $\mathbf{x}$ and $\tilde{\psi}_{\mathbf{x}}$. This correspondence ensures that each dimension of the feature space is directly mapped to a single dimension in the embedding space $\tilde{\psi}_{\mathbf{x}}$.
Our optimization process involves using SGD with positive and negative sampling, similar to recent graph embeddings methods [33, 25]. Positive edge samples are associated with attraction (first term on the right hand side), while negative samples (second term on the right hand side) refer to a pair of nodes that are not connected in the graph, which create repulsion among dissimilar points.
The optimization can be performed jointly for all scales using the 3D tensor. Specifically, our implementation employs an optimization process that operates jointly across all scales, as described in the SGD update rule for iteration $t$ in Eq. (2). At each iteration $t$, associated with either a positive sample (an edge) or a negative sample (a pair of unconnected nodes), the update rule from Eq.(2). is applied. The derivative of the loss function, as detailed in Eq. (4), is computed for all scales $(s_{1},...s_{K})$, with the sign in Eq.(2) depending on whether the sample is positive or negative. Once the optimization concludes, the final embedding is derived from the 3D tensor by summing over the scales, as specified in Eq.(6). Encoding the representation in a 3D tensor offers two key advantages. First, is the computational efficiency: the update rule at each iteration, associated with a pair of nodes, is performed jointly and simultaneously for all scales. This avoids the need for independent optimization per scale, which would significantly increase computational complexity. The second is optimization alignment: jointly optimizing across scales ensures alignment in the optimization process for all scales with respect to a fixed pair of nodes (positive or negative sample). This approach is more effective than optimizing scales independently, as it maintains consistency and improves overall performance.
In this section, we characterize the theoretical properties of the representation power of the SGW operator by considering functions sampled from the Paley-Wiener spaces [27] on combinatorial graphs. Pesenson introduced the Paley-Wiener spaces and demonstrated that Paley-Wiener functions of low type are uniquely determined by their values on certain subsets, known as uniqueness sets $U$. We show that the SGW operator can represent functions $f$ within the Paley-Wiener space more efficiently than the graph Laplacian operator $\mathbf{L}$. This efficiency is demonstrated by showing that the SGW operator is more effective in representing functions with larger bandwidth $\omega$ in the Paley-Wiener spaces (i.e., with higher frequencies) using subsets of nodes from the uniqueness sets $U$.
The $\Lambda$-set is defined as follows: a set of vertices $S\subset V$ is a $\Lambda(S)$-set if any $\varphi\in L_{2}(S)$ satisfies the inequality $||\varphi||\leq\Lambda||\mathbf{L}\varphi||$ for some constant $\Lambda(S)>0$. The infimum of all $\Lambda>0$ for which $S$ is a $\Lambda$-set is denoted by $\Lambda$.
Theorem 1
The proof is provided in the Appendix.
Figure 1: Illustration of the proposed method. (Left) Initial encoding using SGW in (b) and regularized embedding in (c). (Right) Feature embedding space representation, which has a one-to-one correspondence with the input features, can be used along with clustering for interpretability models, such as feature importance.
5 Experimental Results
We demonstrate the advantage of our approach over existing approaches using both synthetic and real datasets. We evaluate our method by testing its clustering performance using the output embedding and comparing it to several representative methods, including UMAP, t-SNE, Diffusion Maps, ISOMAP, PHATE [26], PaCMAP [36], TriMap [1] and HeatGeo [22]. For UMAP, t-SNE, PHATE, PaCMAP, TriMap and HeatGeo, we hyperparameter tune the performance on each dataset, running the method five times for each hyperparameter configuration and taking the average. Then we report the best average score. Additional details are provided in the supplementary material.
5.1 Real Datasets
We study the performance of MS-IMAP compared to other methods for real datasets. We chose a mix of datasets from varied fields: the Census dataset [14] is a financial dataset containing information about individuals extracted from the 1994 US Census; the Zilionis dataset [40] is a biological dataset containing single-cell sequencing data from different types of cells; the Animals with Attributes (AWA) [37] dataset is an image dataset containing images of animals; The Human Activity Recognition (HAR) [2] dataset which consists of sensor data collected from subjects performing six different activities: walking, walking upstairs, walking downstairs, sitting, standing, and lying. Each activity is represented as a feature vector of 561 attributes and 7352 samples; and the Cora dataset [30, 24] is a graph dataset of a citation network consisting of scientific publications. Note that we only use the features without the natural graph associated with Cora. More information about each dataset can be found in the supplementary material. As shown in Table 1, MS-IMAP is competitive and performs the best overall, while also providing an interpretable embedding through its one-to-one correspondence with the original features—unlike the other competing graph embedding methods.
Runtime and Computational Complexity The computational complexity of Multi-Scale IMAP is of $O(NDlog(N))$. The execution time of our method with Python code implementation with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM on the Cancer QC data-set of 48,969 samples and 306 features took 27 mins, on Cora 5 min, on the AWA dataset 1.35 mins and on Census 7 min. The increase in running times compared to methods such as UMAP is primarily due to our extension of manifold learning to include interpretability. Our approach generally requires the embedding dimensionality to be at least equal to the input dimensionality of the features, which adds computational complexity but ensures meaningful feature representation and interpretation. Our method is practical for handling datasets with millions of points and dozens to hundreds of features within a few hours. In fields such as finance and genomics/biology, where typical dataset sizes range from under 100,000 to several million data points, analysis is often conducted offline once the graph embedding is computed. Runtimes of up to a few hours are generally acceptable and do not present a significant challenge, making the method well-suited for real-world applications. Additional details are provided in the Appendix.
Data HAR Census Zilionis AWA Cora
Method / Accuracy ARI AMI ARI AMI ARI AMI ARI AMI ARI AMI
Features 0.45 0.58 0.11 0.15 0.48 0.63 0.07 0.16 0.01 0.07
ISOMAP 0.55 0.65 0.18 0.09 0.44 0.56 0.52 0.60 0.12 0.17
Diffusion Maps N/A N/A 0.06 0.16 0.01 0.02 0.22 0.42 0.03 0.07
UMAP 0.61 0.73 0.23 0.15 0.52 0.74 0.74 0.82 0.33 0.36
t-SNE 0.71 0.77 0.17 0.17 0.41 0.70 0.73 0.80 0.21 0.27
PHATE 0.55 0.66 0.20 0.09 0.65 0.72 0.66 0.73 0.12 0.22
TriMap 0.59 0.72 0.2 0.14 0.62 0.77 0.74 0.81 0.15 0.2
PaCMAP 0.62 0.75 0.24 0.15 0.55 0.76 0.75 0.81 0.21 0.26
HeatGeo 0.6 0.69 0.15 0.10 N/A N/A 0.65 0.74 0.17 0.20
MS-IMAP 0.67 0.8 0.23 0.15 0.70 0.76 0.74 0.81 0.33 0.36
Table 1: Clustering results comparison using ARI and AMI on the Census, Zilionis, AWA, and Cora datasets. Note HeatGeo had code execution issues on the Zilionis dataset.
6 Feature Importance using Explicit Correspondence with Embedding Features
The correspondence established between the features and the embedding space offers a valuable tool for assessing feature importance. We explore two alternative methods for measuring feature importance, leveraging the constructed embedding space, where each dimension can be traced back to its corresponding feature in the original space. Specifically, we test these methods in an unsupervised or self-supervised setting where the subset of features are selected without using label information.
6.1 Assessing Embedding Feature Importance using the Laplacian Score
The first method utilizes the Laplacian Score (LS) [19], which assesses the importance of each feature by examining its correlation with the eigenvalues of the graph Laplacian. The significance of the variable corresponding to the $l$-th coordinate, represented by the feature $f_{l}\in\mathbb{R}^{N}$, is inferred from the relevance of its corresponding coordinate $l$ in the embedding space, $\tilde{\psi}_{\mathbf{x}_{l}}$. The Laplacian Score is calculated with respect to the embedding features $\tilde{\psi}_{\mathbf{x}}$ using the graph Laplacian $\mathbf{L}$ and the degree matrix $\mathbf{D}$. To compute the Laplacian Score, we first subtract the mean and then calculate it as follows: $L_{s}(\tilde{\psi}_{\mathbf{x}})_{l}=\frac{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{L}(\tilde{\psi}_{\mathbf{x}})_{l}}{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{D}(\tilde{\psi}_{\mathbf{x}})_{l}}$
Smaller scores indicate that the embedding feature $(\tilde{\psi}_{\mathbf{x}})_{l}$ has a greater projection onto the subspace of eigenvectors corresponding to the smallest eigenvalues, signifying higher importance concerning the global graph structure. Therefore, the feature importance of $(\tilde{\psi}_{\mathbf{x}})_{l}$ can be directly interpreted as the importance of the corresponding original feature.
6.2 Assessing Importance of Embedding Features using Mutual Information
In the second method, we propose a self-supervised approach that uses Mutual Information (MI) to estimate the importance of each dimension in the embedding space, which can then be explicitly linked to the original features. First, we apply $k$-means to cluster the MS-IMAP embedding space and use the resulting clusters as pseudo-labels. We then compute MI to quantify the shared information between features and these pseudo-labels, similar to how MI would be computed with ground truth labels. This step allows MI to measure how much each feature contributes to the cluster assignments, highlighting those most informative in distinguishing clusters. The MI-Based Feature Importance identify features that are informative about cluster assignments. Thus, masking these features would especially degrade the clustering performance. We summarize the MI-based unsupervised feature selection in Algorithm 2. In comparison to the MI based feature importance method, using the Laplacian score [19] we have that the embedding may have combined original features in a way that makes individual Laplacian Scores less indicative of feature importance. The Laplacian score prioritize neighborhood preservation, which might not directly translate to cluster discrimination efficacy. On the other hand, while the proposed MI based feature importance identifies features that are directly informative for distinguishing clusters, one of its possible limitations may be relaying on the quality of pseudo-labels, which may not align with true clusters.
  Input: MS-IMAP embeddings $\psi_{\mathbf{x}}$, number of clusters $k$
1  Output: MI scores for each feature
  Step 1: Apply k-means clustering on the embedding space $\psi_{\mathbf{x}}$ to obtain cluster labels
[…]
  Step 2: Use the resulting cluster labels as pseudo-labels
[…]
  Step 3: Initialize an empty list to store mutual information (MI) values for each feature
[…]
  Step 4: For each feature $i$ in the original data or embedding space:
            Step 4.1: Compute mutual information between feature $\psi_{\mathbf{x}}(i)$ and the pseudo-labels
4            Step 4.2: $MI\_score=\text{MI}(\psi_{\mathbf{x}}(i),c({\psi_{\mathbf{x}}})$)
5            Step 4.3: Append $MI\_score$ to mi_scores
  Step 5: Output the list of MI scores
            Return mi_scores
Algorithm 2 MI-based Unsupervised Feature Selection
Figure 2: Clustering results on Cora using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
6.3 Assessing Feature Importance via Embedding Correspondence
We demonstrate the application of MS-IMAP for feature importance by evaluating clustering performance using $k$-means as a function of the number of features, based on the most significant features identified by two methods: the Laplacian score and Mutual Information-based feature selection. We compared the clustering results (using ARI and AMI) when $k$-means is applied to different subsets of embedding features selected by the Laplacian score and MI-based methods, with results obtained by applying the same feature selection methods directly to the original feature space.
Unlike most manifold learning and graph embedding methods, MS-IMAP produces embeddings where the dimensions can be directly mapped to the original feature space, enabling the assessment of individual feature contributions.
On the Cora dataset, Figure 2 shows that for both feature selection methods, the clustering accuracy obtained using the MS-IMAP embedding is significantly higher than when using the raw features across different subsets of features. This improvement is due to the robustness of the MS-IMAP embedding, while still offering interpretability by linking the embedding dimensions to the original features.
7 Discussion
Identifying the key factors of high-dimensional datasets is a necessary ingredient to take advantage of unlabelled data in many practical applications. In this article, we introduce a novel contrastive learning framework for manifold learning via graph embeddings, capitalizing on both low and high-frequency information. In particular, we use SGW to construct a multi-scale graph representation of the underlying input feature space. We then use an SGD-based optimization scheme together with innovative strategies such as 3D Tensor encoding to derive the embeddings. To motivate the use of the spectral graph wavelet (SGW) for representation, we examine the theoretical properties of the SGW by studying functions in Paley-Wiener spaces on combinatorial graphs. We prove that the SGW operator provides a more effective representation through the concept of the $\Lambda$-set. Finally, we demonstrate that the embeddings are interpretable using two alternative methods for unsupervised and self-supervised feature importance, applied to both the embedding and original feature spaces. The construction of our methodology inherently links the original and transformed feature spaces, a capability that is notably absent in most existing non-linear manifold learning techniques, to the best of our knowledge.
Limitations and future work include the following. Firstly, we rely on the assumption that the features, and thus the similarity measure, adequately capture information to reveal geodesic distances between manifold points. Secondly, utilizing the current optimization framework with SGD may present difficulties when extending to out-of-sample problems. Furthermore, although our embedding feature correspondence proves valuable in assessing feature importance through a global method, there remains potential for enhancement through the incorporation of local methods.
Appendix
8 Encoding Method 1
[…]
denoting the concatenation of the vectors corresponding to the multi-scale representation $\psi_{{f}_{i}}(s_{k},:)$ and $\psi_{{f}_{j}}(s_{k},:)$. For method 1 all features and all scales are concatenated together, the resulting matrix $\psi_{\mathbf{x}}$ can be represented as
where $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$.
@@ 1. Optimize Embedding Method 1:
$\frac{\partial\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})}{\partial\psi_{\mathbf{x}_{i}}}\sim\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}}(r_{ij})-\sum_{j}\frac{1}{||r_{ij}||^{2}}v_{ij}^{\psi_{\mathbf{x}}}(\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}})$ (8)
where $r_{ij}=\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}}$, and $\psi_{\mathbf{x}_{i}}$ corresponds to the vector associated with node $i$ multi-scale representation concatenated for all the graph signals and all spectral bands. since for method 1 all features and all scales are concatenated together,
9 Additional experiments: Embedding Feature Importance using Mutual Information and Laplacian Score
We include additional experiments that demonstrate the application of MS-IMAP for feature importance by evaluating clustering performance using $k$-means as a function of the number of features (in addtion to the one found in the Section 6), based on the most significant features identified by the Laplacian score and Mutual Information-based feature selection. We compared the clustering results (using ARI and AMI) when $k$-means is applied to different subsets of embedding features selected by the Laplacian score and MI-based methods, with results obtained by applying the same feature selection methods directly to the original feature space.
On the Zillions dataset, Figures 5 and 6 shows that for both feature selection methods, the clustering accuracy obtained using the MS-IMAP embedding is significantly higher than when using the raw features across different subsets of features. This improvement is due to the robustness of the MS-IMAP embedding, while still offering interpretability by linking the embedding dimensions to the original features. We also show in Figure 4 the clustering accuracy obtained using the MS-IMAP copared to using the raw features using AMI, demonstrating similar observations regarding the advantages of MS-IMAP.
Figure 4: Clustering results (AMI) on Cora using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
Figure 5: Clustering results (ARI) on Zillions using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
Figure 6: Clustering results (AMI) on Zillions using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
10 Dataset Details
In Figure 7, we show example clusterings for each of the methods mentioned in the Experimental Results section of the main paper.
Cora. The Cora dataset has 2,708 scientific publications categorized into seven classes. The network has 5,429 links and each publication is described as a binary word vector indicating the presence or absence of a word. The dictionary consists of 1,433 words. In our experiments, we do not use the given graph, and only use the features. [30, 24].
HAR. The Human Activity Recognition (HAR) dataset consists of sensor data collected from 30 subjects performing six different activities: walking, walking upstairs, walking downstairs, sitting, standing, and lying. The data was recorded using accelerometer and gyroscope sensors embedded in a smartphone worn on the waist. Each activity is represented as a feature vector of 561 attributes and 7352 samples, extracted from the raw sensor signals through signal processing techniques. In our experiments, we do not use the temporal dependencies and only utilize the extracted features.
Figure 7: Clustering results using two moons
11 Experimental Results on two moons Synthetic Dataset